# MOFA-Video: Controllable Image Animation via Generative Motion Field Adaptions in Frozen Image-to-Video Diffusion Model

### v1

@@ Abstract
We present MOFA-Video, an advanced controllable image animation method that generates video from the given image using various additional controllable signals (such as human landmarks reference, manual trajectories, and another even provided video) or their combinations. This is different from previous methods which only can work on a specific motion domain or show weak control abilities with diffusion prior. To achieve our goal, we design several domain-aware motion field adapters (i.e., MOFA-Adapters) to control the generated motions in the video generation pipeline. For MOFA-Adapters, we consider the temporal motion consistency of the video and generate the dense motion flow from the given sparse control conditions first, and then, the multi-scale features of the given image are wrapped as a guided feature for stable video diffusion generation. We naively train two motion adapters for the manual trajectories and the human landmarks individually since they both contain sparse information about the control. After training, the MOFA-Adapters in different domains can also work together for more controllable video generation.
11footnotetext: Corresponding author
@@ 1 Introduction
With the development of digital devices, current methods attempt to animate the image by computer vision algorithms [24, 32, 36, 17, 8, 27, 21, 16, 30, 40, 50, 34, 13]. However, it faces several limitations. On the one hand, these methods often focus on limited categories of animated objects such as fluids [16, 24, 25], human hairs [37], and human body/face [50, 6, 9, 17, 32, 36, 34, 8, 13, 27]. Thanks to the domain knowledge of each specific type, these methods often have fully controllable abilities of the scenes. e.g., SadTalker [50] can produce the accurate human face animation by audio and the given face. Text2Cinemagraph [25] produces the natural animation of the water using text description. For control abilities, these methods usually follow the rule of learning the video via self-supervised decomposition and then animating via new driving signals. However, due to the limitation of natural animation prior, these methods fail in the general image space because of the diversity of the general domain knowledge.
Unlike previous in-domain image animation, current diffusion-based Image-to-Video (I2V) methods learn to generate the video from the image in an end-to-end fashion. Thanks to the large-scale generative prior of the text-to-image model, i.e., Stable Diffusion [29], these methods [7, 39, 11, 2, 1] have proved the possibility of open domain image animation. However, their generated contents might be different with the given image [1, 2, 11, 39], and often generate simple motions by text descriptions [39, 1, 2] or just simple idle animation [7]. These drawbacks limit their applications for real-world image animation tasks, where users often need to create more controllable videos as in previous in-domain image animation algorithms.
@@
To this end, we present MOFA-Video to add the additional different motion control abilities to the generatic video diffusion model (Stable Video Diffusion [7] in our case), inspired by ControlNet [48]. In detail, to animate an input image into a video according to the sparse control signals from multiple domains, we design a novel MOFA-Adapter that serves as an additional adapter on the pretrained video diffusion model so that the motions of the video can be controlled. Different from the previous ControlNet-like Adapter [38, 51] for video generation, the proposed MOFA-Adapter models the frame-wise motion explicitly. In detail, we first utilize the given sparse motion hint to generate the dense motion fields using a sparse-to-dense motion generation network, then, we warp the multi-scale features of the first frame as the conditional features for the diffusion generation process. This sparse-to-dense motion generation provides a good balance between the provided motion guidance and the generation process, providing high-quality animation results with good temporal consistency. We also think about the problems of there containing multiple motion domains. Thus, we train multiple MOFA-Adapters by considering these tasks as sparse control point generation problems, including the open-world manual trajectories, human facial animations, etc. Besides, since the parameters of the video diffusion model are fixed, we can perform the motion control abilities across multiple domains jointly, e.g., the human face and background objects and the camera movement jointly. We give more detailed applications and examples in the experiments.
@@ 2.1 Controllable Image Animation
Image animation has a long story in computer vision and graphics. Since it is a very ill-posed problem, previous work only focuses on the specific domain. e.g., previous methods to generate the video from the image in an unsupervised and stochastic manner [41, 20, 16, 22] using generative model, e.g., generative adversarial networks (GAN). These methods do not provide the control handle of the generation, limiting its applications on real-world applications. On the other hand, controlling the motion of the generation is also difficult for in-the-wild images. To this end, recent works aim to control the fluids [24, 25], human pose/face [9, 17, 32, 36, 34, 8, 13, 27] only, other than the general scenes. Different from these methods, we propose a novel and unified framework to model the motions from different domains and make it work singly and together in a pre-trained Video Diffusion Model.
@@
In the realm of downstream tasks utilizing VDMs for video-related applications, there exists a category of work known as Image-to-Video Diffusion Models (I2Vs) [1, 2, 4, 39, 3, 7]. The goals of these models are similar to image animation, but they differ in some aspects. The primary difference is that most of these methods employ an image encoder to extract semantic features from a reference image to guide video generation, without requiring the generated video to precisely include the given image as the first frame. However, current I2V models only generate the video from the given image and aim to control the motions from the text description, we argue it is somewhat hard for text-based motion generation since they contain limited motion priors. Some related methods also try to control the generation of the motion. e.g., DragNUWA [45] model the trajectory of the generation via adaptive normalization, which shows weak spatial correction. MotionCtrl [35] tries to control the object and camera motions of the T2V model, however, it is hard for the generation process since there is no world coordinate system for text-to-video generation. The most related controllable image animation is the concurrent work, MotionI2V [31]. However, it only focuses on the natural motion of the objects. which is hard to fully control the motions from other domains, e.g., the human face or body. Differently, we aim to create a method which is built on a pre-trained base model, i.e., SVD, and also enjoys the motion hints from different domains using different adapters.
@@ 3 Method
We aim to generate videos from the given reference image and additional motion control signals in multiple motion domains (e.g., hand-crafted trajectories, human landmark sequences, dense motion flows, etc.) under a unified framework, so they can share a unified network structure and jointly work together like Multi-ControlNet [48]. To achieve this goal, as shown in Fig. 2, we design a generative motion field adapter (MOFA-Adapter) that can accept sparse motion control signals as the condition and produce the detailed control abilities for a frozen Stable Video Diffusion [7] Model. We train the proposed MOFA-Adapter in two different motion domains individually and provide various applications based on each model and their combinations.
Below, we first introduce the structure of the proposed MOFA-Adapter in Sec. 3.1. Then, we give the details of how we train the domain-aware MOFA-Adapter for the video diffusion model in Sec. 3.2. Finally, in Sec. 3.3, we give the inference details of the proposed method and various additional applications.
@@ 3.1 Generative Motion Field Adapter (MOFA-Adapter)
As shown in Fig. 3, the proposed adapter is based on a reference encoder, a sparse-to-dense (S2D) motion generator for sampled motion hints, and the feature fusion encoder to add the warped feature back to the pretrained video diffusion model. In detail, the reference image encoder is a multi-scale convolutional feature encoder, which extracts the multi-scale features of the first frame for warping, with each stage built by Conv-SiLU-ZeroConv [48] as the basic block. As for the sparse-to-dense motion generator, we use the same network structure as CMP [47] for adaption. This network is also a convolutional neural network which accepts the first frame image and the sparse hint of the motion, and produces the dense motion fields. More details of the network structure can be founded in the original paper [47]. When the dense motion field is generated, we warp the referenced features and then add them to the feature map of the corresponding levels of a copied SVD encoder, which is then added to the feature space of the decoder of the pre-trained SVD similar to ControlNet [48].
@@ 3.2 Training MOFA-Adapters on Stable Video Diffusion
Fig. 2 illustrates the training pipeline of our whole framework. Given a video clip $V\in\mathbb{R}^{L\times 3\times H\times W}$ in $L$ frames, we first extract the sparse motion vectors which serves as the input of the S2D network. e.g., for open-domain, we handle the motion hint as the sparse motion vectors sampled from the extracted dense optical flow. As for human motion, we generate the motion hints from structural key-points, e.g., facial landmarks. Below, we give the details of each specific type:
Sparse Motion Vectors from Dense Optical Flow. By considering the dense optical flow as a general motion representation between video frames, we first utilize Unimatch [42] to extract the forward optical flow as $F\in\mathbb{R}^{(L-1)\times 2\times H\times W}$. The flow sequences between the $i$-th and $(i+1)$-th frames are represented as $F_{i}\in\mathbb{R}^{2\times H\times W}$. Based on the flow sequence $F$, we sample $n$ spatial points from each frame $F_{i}$ using the watershed sampling strategy [47], resulting in a sparse mask $M^{s}\in\mathbb{R}^{H\times W}$, where the value for the sampled spatial points is set to 1, and other points are set to 0. We then calculate the sparse motion vectors $F^{s}\in\mathbb{R}^{(L-1)\times 2\times H\times W}$ as:
@@ 3.3 Inferences
After training, our method can generate video from a single image and the given control signals, e.g., handcrafted trajectories, facial key-points, etc. In the following parts, we introduce several inference pipelines and applications of our method, including hand-craft trajectory-based animation, motion brush masking, portrait animation with key-points, etc.
@@
Hand-Crafted Trajectories for Objects/Camera Movement. Given an input image $I\in\mathbb{R}^{H\times W\times 3}$, the users can draw multiple motion trajectories on the image, each of which can be represented as: $\mathcal{T}=[(x_{0},y_{0}),(x_{P-1},y_{P-1})]$. Take one trajectory for example, we bicubicly interpolate this trajectory to $L$ points, i.e., $\hat{\mathcal{T}}\in\mathbb{R}^{L\times 2}$ and calculate the sparse motion hints $F^{s}\in\mathbb{R}^{(L-1)\times 2\times H\times W}$ between the frames as:
@@
As a result, we can create several trajectories to indicate the motions of the specific object individually. As shown in Fig. 4, we can perform drags in both linear and non-linear fashion since our model is able to handle both these two types of motion via optical flows. Besides, by adding relatively dense guidance on the whole images, our method also capability of the motions of the whole scene to perform as “camera motion”.
Regional Motion Brushes. Besides simple hand-craft trajectories, we can generate the regional motion animation based on the trajectory model. This feature is different from current implicitly drag-based methods (DragNUWA [45]), where the animated region is determined by the model itself instead of the user. Specifically, the user could additionally provide a binary motion mask $M\in\mathbb{R}^{H\times W}$ to control the animation region. Given this mask, we divide the trajectory set $\mathcal{T}$ into two groups: $\mathcal{T}_{\text{in}}$ and $\mathcal{T}_{\text{out}}$. For each trajectory, we distribute it to $\hat{\mathcal{T}}_{\text{in}}$ if its start points lie in the mask, otherwise to $\hat{\mathcal{T}}_{\text{out}}$. We then respectively obtain the dense optical flow $F_{in}$ and $F_{out}$ via the S2D network, and mask out all the values that is located outside the mask. We then obtain the final dense optical flow by combining the masked $F_{in}$ and $F_{out}$. We demonstrate the corresponding results in Fig. 4.
@@
3.3.1 Portrait Image Animation.
Besides natural object movement, we also consider the motion domain of the human, specifically the simple face movement, which is more detailed and structural. For inference animation, we first use an audio-driven talking face generation method, i.e., SadTalker [50], to produce the 3D facial landmark from a single image, then, as introduced in Sec. 3.1, we generate the sparse motion hints and use the trained MOFA-Adapter for motion generation. Fig. 5 demonstrates image animation results via audio-driven facial key-points generation [50] and a single image. Notice that, since we use facial landmarks as the intermediate representation, we can also drive the animation with another facial video or manual facial blendshapes animation.
3.3.2 Longer Animation.
Despite the significant generation ability provided by the SVD, the frame length of its generated video is extremely limited to the specific frames (e.g., 14 in our case). The computational complexity of temporal attention scales quadratically with the number of frames, making it difficult to directly generate ultra-long videos. On the other hand, ensuring consistency becomes more challenging with the increase in the number of frames. To solve this limitation, we propose the periodic sampling strategy to resolve the frame number issue and generate longer videos, inspired by Gen-L-Video [33]. As illustrated in Fig. 6, in each diffusion step, we periodically sample 14 frames as one group and denoise them using our SVD model. Each latent group has an overlapping of 7 frames to provide temporal smoothness. After all the groups are sampled, we average the predicted noise for each frame and obtain the final latent for this diffusion step.
@@
3.3.3 Plugin Combinations.
Our careful design enables us to animate the given image by simultaneously using different control signals from multiple modalities without any retraining. This is achieved by combining multiple MOFA-Adapters for the animation similar to Multi-ControlNet [48]. In Fig. 7, we show an example of controlling the facial expression of one portrait while animating other objects such as backgrounds via hand-craft trajectories. Please refer to the supplementary materials for more experiment results and technical details.
@@ 4 Experiment
4.0.1 Implementation Details.
For training, we train the proposed trajectories-based MOFA-Adapter on WebVid-10M [5] using 8 NVIDIA A100 GPU. The base model is Stable Video Diffusion [7]. We use AdamW [23] as an optimizer. During training, we randomly sample 14 video frames with a stride of 4. The learning rate is set to $2\times 10^{-5}$ with a resolution of $256\times 256$. We first train the model as a flow-based reconstruction model by removing the S2D motion generator and directly taking the first frame together with the estimated optical flow from Unimatch [42]. After that, we add the S2D network and finetune the whole adapter. For portrait-based MOFA-Adapter, we train our model on a self-collected human video dataset. The pre-trained DWPose [43] is used to extract the facial landmark and other training settings are the same as the trajectories-based model. Please refer to the supplementary for more details.
4.0.2 Evaluations Metrics.
For the trajectory-based MOFA-Adapter, we use 1000 samples from the test set of WebVid and respectively calculate LPIPS [49], FID [14], and FVD [46]. We also calculate the cosine similarity of the CLIP [28] embedding of the consecutive generated frames to evaluate the frame consistency of the results. To evaluate the keypoint-based MOFA-Adapter, we choose 40 generated video examples, each of which contains 196 frames, and calculate the following two metrics inspired by previous work [50]: 1) Cumulative Probability Blur Detection (CPBD) [26] to evaluate the sharpness of the generated frames, and 2) identity embedding from ArcFace [12] between the source images and the generated frames to evaluate the fidelity of the generated results. Apart from the quantitative metrics, we also conduct user studies to perceptually evaluate the result of different methods.

### v3

@@ Abstract
We present MOFA-Video, an advanced controllable image animation method that generates video from the given image using various additional controllable signals (such as human landmarks reference, manual trajectories, and another even provided video) or their combinations. This is different from previous methods which only can work on a specific motion domain or show weak control abilities with diffusion prior. To achieve our goal, we design several domain-aware motion field adapters (i.e., MOFA-Adapters) to control the generated motions in the video generation pipeline. For MOFA-Adapters, we consider the temporal motion consistency of the video and generate the dense motion flow from the given sparse control conditions first, and then, the multi-scale features of the given image are wrapped as a guided feature for stable video diffusion generation. We naively train two motion adapters for the manual trajectories and the human landmarks individually since they both contain sparse information about the control. After training, the MOFA-Adapters in different domains can also work together for more controllable video generation. Codes available: https://github.com/MyNiuuu/MOFA-Video
11footnotetext: Corresponding authors
@@ 1 Introduction
With the development of digital devices, current methods attempt to animate the image using computer vision algorithms [24, 32, 36, 17, 8, 27, 21, 16, 30, 40, 50, 34, 13]. However, it faces several limitations. On the one hand, these methods often focus on limited categories of animated objects such as fluids [16, 24, 25], human hairs [37], and human body/face [50, 6, 9, 17, 32, 36, 34, 8, 13, 27]. Thanks to the domain knowledge of each specific type, these methods often have fully controllable abilities of the scenes. e.g., SadTalker [50] can produce the accurate human face animation by audio and the given face. Text2Cinemagraph [25] produces the natural animation of water using text description. For control abilities, these methods usually follow the rule of learning the video via self-supervised decomposition and then animating via new driving signals. However, due to the limitation of natural animation prior, these methods fail in the general image space because of the diversity of the general domain knowledge.
Unlike previous in-domain image animation, current diffusion-based Image-to-Video (I2V) methods learn to generate the video from the image in an end-to-end fashion. Thanks to the large-scale generative prior of the text-to-image model, i.e., Stable Diffusion [29], these methods [7, 39, 11, 2, 1] have proved the possibility of open-domain image animation. However, their generated contents might be different from the given image [1, 2, 11, 39], and often generate simple motions by text descriptions [39, 1, 2] or just simple idle animation [7]. These drawbacks limit their applications for real-world image animation tasks, where users often need to create more controllable videos as in previous in-domain image animation algorithms.
@@
To this end, we present MOFA-Video to add the additional different motion control abilities to the generatic video diffusion model (Stable Video Diffusion [7] in our case), inspired by ControlNet [48]. In detail, to animate an input image into a video according to the sparse control signals from multiple domains, we design a novel MOFA-Adapter that serves as an additional adapter on the pre-trained video diffusion model so that the motions of the video can be controlled. Different from the previous ControlNet-like Adapter [38, 51] for video generation, the proposed MOFA-Adapter models the frame-wise motion explicitly. In detail, we first utilize the given sparse motion hint to generate the dense motion fields using a sparse-to-dense motion generation network, then, we warp the multi-scale features of the first frame as the conditional features for the diffusion generation process. This sparse-to-dense motion generation provides a good balance between the provided motion guidance and the generation process, providing high-quality animation results with good temporal consistency. We also think about the problems of there containing multiple motion domains. Thus, we train multiple MOFA-Adapters by considering these tasks as sparse control point generation problems, including open-world manual trajectories, human facial animations, etc. In addition, since the parameters of the video diffusion model are fixed, we can jointly perform motion control abilities across multiple domains, e.g., the human face and background objects and the camera movement. We give more detailed applications and examples in the experiments.
@@ 2.1 Controllable Image Animation
Image animation has a long history in computer vision and graphics. Since it is a very ill-posed problem, previous work only focuses on the specific domain. e.g., previous methods to generate the video from the image in an unsupervised and stochastic manner [41, 20, 16, 22] using generative model, e.g., generative adversarial networks (GAN). These methods do not provide the control handle of the generation, limiting its applications on real-world applications. On the other hand, controlling the motion of the generation is also difficult for in-the-wild images. To this end, recent works aim to control the fluids [24, 25], human pose/face [9, 17, 32, 36, 34, 8, 13, 27] only, other than the general scenes. Different from these methods, we propose a novel and unified framework to model the motions from different domains and make it work singly and together in a pre-trained Video Diffusion Model.
@@
In the realm of downstream tasks utilizing VDMs for video-related applications, there exists a category of work known as Image-to-Video Diffusion Models (I2Vs) [1, 2, 4, 39, 3, 7, 52]. However, current I2V models only generate the video from the given image and aim to control the motions from the text description, we argue it is somewhat hard for text-based motion generation since they contain limited motion priors. Some related methods also try to control the generation of the motion. e.g., DragNUWA [45] model the trajectory of the generation via adaptive normalization, which shows weak spatial correction. MotionCtrl [35] tries to control the object and camera motions of the T2V model, however, it is hard for the generation process since there is no world coordinate system for text-to-video generation. The most related controllable image animation is the concurrent work MotionI2V [31]. However, it focuses only on the natural motion of the objects. which is hard to fully control the motions from other domains, e.g., the human face or body. Differently, we aim to create a method which is built on a pre-trained base model, i.e., SVD, and also enjoys the motion hints from different domains using different adapters.
@@ 3 Method
We aim to generate videos from the given reference image and additional motion control signals in multiple motion domains (e.g., hand-crafted trajectories, human landmark sequences, dense motion flows, etc.) in a unified framework, so they can share a unified network structure and jointly work together like Multi-ControlNet [48]. To achieve this goal, as shown in Fig. 2, we design a generative motion field adapter (MOFA-Adapter) that can accept sparse motion control signals as the condition and produce detailed control abilities for a frozen Stable Video Diffusion [7] Model. We train the proposed MOFA-Adapter in two different motion domains individually and provide various applications based on each model and their combinations.
In the following, we first introduce the structure of the proposed MOFA-Adapter in Sec. 3.1. Then, we give details of how we train the domain-aware MOFA-Adapter for the video diffusion model in Sec. 3.2. Finally, in Sec. 3.3, we give the inference details of the proposed method and various additional applications.
@@ 3.1 Generative Motion Field Adapter (MOFA-Adapter)
As shown in Fig. 3, the proposed adapter is based on a reference encoder, a sparse-to-dense (S2D) motion generator for sampled motion hints, and the feature fusion encoder to add the warped feature back to the pre-trained video diffusion model. In detail, the reference image encoder is a multi-scale convolutional feature encoder, which extracts the multi-scale features of the first frame for warping, with each stage built by Conv-SiLU-ZeroConv [48] as the basic block. For the sparse-to-dense motion generator, we use the same network structure as CMP [47] for adaption. This network is also a convolutional neural network which accepts the first frame image and the sparse hint of the motion, and produces the dense motion fields. More details of the network structure can be founded in the original paper [47]. When the dense motion field is generated, we warp the referenced features and then add them to the feature map of the corresponding levels of a copied SVD encoder, which is then added to the feature space of the decoder of the pre-trained SVD similar to ControlNet [48].
@@ 3.2 Training MOFA-Adapters on Stable Video Diffusion
Fig. 2 illustrates the training pipeline of our whole framework. Given a video clip $V\in\mathbb{R}^{L\times 3\times H\times W}$ in $L$ frames, we first extract the sparse motion vectors which serve as input to the S2D network. e.g., for open-domain, we handle the motion hint as the sparse motion vectors sampled from the extracted dense optical flow. As for human motion, we generate the motion hints from structural key-points, e.g., facial landmarks. In the following, we give the details of each specific type:
Sparse Motion Vectors from Dense Optical Flow. By considering the dense optical flow as a general motion representation between video frames, we first utilize Unimatch [42] to extract the forward flow as $F\in\mathbb{R}^{(L-1)\times 2\times H\times W}$, where $F_{i}\in\mathbb{R}^{2\times H\times W}$ represents the optical flow from the 00-th to the $(i+1)$-th frames. Based on the flow sequence $F$, we sample $n$ spatial points for each frame $F_{i}$ using the watershed sampling strategy [47]. Specifically, we first obtain a sparse mask $M^{s}\in\mathbb{R}^{H\times W}$, where the value for the sampled spatial points is set to 1, and other points are set to 0. We then calculate the sparse motion vectors $F^{s}\in\mathbb{R}^{(L-1)\times 2\times H\times W}$ as:
@@ 3.3 Inferences
After training, our method can generate video from a single image and the given control signals, e.g., handcrafted trajectories, facial key-points, etc. In the following parts, we introduce several inference pipelines and applications of our method, including hand-crafted trajectory-based animation, motion brush masking, and portrait animation with key-points, etc.
@@
Hand-Crafted Trajectories for Objects/Camera Movement. Given an input image $I\in\mathbb{R}^{H\times W\times 3}$, the users can draw multiple motion trajectories on the image, each of which can be represented as $P$ joint points $\mathcal{T}\in\mathbb{R}^{P\times 2}=[(x_{0},y_{0}),(x_{1},y_{1}),...,(x_{P-1},y_{P-1})]$. Take one trajectory for example, we bicubicly interpolate this trajectory to $L$ points, i.e., $\hat{\mathcal{T}}\in\mathbb{R}^{L\times 2}$ and calculate the sparse motion hints $F^{s}\in\mathbb{R}^{(L-1)\times 2\times H\times W}$ between the frames as:
@@
As a result, we can create several trajectories to indicate the motions of the specific object individually. As shown in Fig. 4, we can handle trajectories in both linear and non-linear fashion since our model is able to handle both these two types of motion via optical flows. In addition, by adding relatively dense guidance on the whole images, our method also has the capability to control the motions of the entire scene to perform as “camera motion”.
Regional Motion Brushes. Besides simple hand-craft trajectories, we can generate the regional motion animation based on the trajectory model. This feature is different from current implicitly drag-based methods (DragNUWA [45]), where the animated region is determined by the model itself instead of the user. Specifically, the user could additionally provide a binary motion mask $M\in\mathbb{R}^{H\times W}$ to control the animation region. Given this mask, we divide the trajectory set $\mathcal{T}$ into two groups: $\mathcal{T}_{\text{in}}$ and $\mathcal{T}_{\text{out}}$. For each trajectory, we distribute it to $\hat{\mathcal{T}}_{\text{in}}$ if its start points lie in the mask, otherwise to $\hat{\mathcal{T}}_{\text{out}}$. We then respectively obtain the dense optical flow $F_{in}$ and $F_{out}$ via the S2D network, and mask out all the values that are located outside the mask. We then obtain the final dense optical flow by combining the masked $F_{in}$ and $F_{out}$. We demonstrate the corresponding results in Fig. 4.
@@
Portrait Image Animation. Besides natural object movement, we also consider the motion domain of the human, specifically the simple face movement, which is more detailed and structural. For inference animation, we first use an audio-driven talking face generation method, i.e., SadTalker [50], to produce facial landmarks from a single image, then, as introduced in Sec. 3.1, we generate the sparse motion hints and use the trained MOFA-Adapter for motion generation. Fig. 5 demonstrates image animation results via audio-driven facial key-points generation [50] and a single image. Notice that, since we use facial landmarks as the intermediate representation, we can also drive the animation with another facial video or manual facial blendshapes animation.
Longer Animation. Despite the significant generation ability provided by the SVD, the frame length of its generated video is extremely limited to the specific frames (e.g., 14 in our case). The computational complexity of temporal attention scales quadratically with the number of frames, making it difficult to directly generate ultra-long videos. On the other hand, ensuring consistency becomes more challenging with the increase in the number of frames. To solve this limitation, we propose the periodic sampling strategy to resolve the frame number issue and generate longer videos, inspired by Gen-L-Video [33]. As illustrated in Fig. 6, in each diffusion step, we periodically sample 14 frames as one group and denoise them using our SVD model. Each latent group has an overlapping of 7 frames to provide temporal smoothness. After all the groups are sampled, we average the predicted noise for each frame and obtain the final latent for this diffusion step.
@@
Plugin Combinations. Our careful design enables us to animate the given image by simultaneously using different control signals from multiple modalities without any retraining. This is achieved by combining multiple MOFA-Adapters for the animation similar to Multi-ControlNet [48]. In Fig. 7, we show an example of controlling the facial expression of one portrait while animating other objects such as backgrounds via hand-craft trajectories. Please refer to the supplementary materials for more experiment results and technical details.
@@ 4 Experiment
Implementation Details. For training, we train the proposed trajectories-based MOFA-Adapter on WebVid-10M [5] using 8 NVIDIA A100 GPU. The base model is Stable Video Diffusion [7]. We use AdamW [23] as an optimizer. During training, we randomly sample 14 video frames with a stride of 4. The learning rate is set to $2\times 10^{-5}$ with a resolution of $256\times 256$. We first train the model as a flow-based reconstruction model by removing the S2D motion generator and directly taking the first frame together with the estimated optical flow from Unimatch [42]. After that, we add the S2D network and finetune the whole adapter. For portrait-based MOFA-Adapter, we train our model on a self-collected human video dataset. The pre-trained DWPose [43] is used to extract the facial landmark and other training settings are the same as the trajectories-based model. Please refer to the supplementary for more details.
Evaluation Metrics. For the trajectory-based MOFA-Adapter, we use 1000 samples from the test set of WebVid and respectively calculate LPIPS [49], FID [14], and FVD [46]. We also calculate the cosine similarity of the CLIP [28] embedding of the consecutive generated frames to evaluate the frame consistency of the results. To evaluate the keypoint-based MOFA-Adapter, we choose 40 generated video examples, each of which contains 196 frames, and calculate the following two metrics inspired by previous work [50]: 1) Cumulative Probability Blur Detection (CPBD) [26] to evaluate the sharpness of the generated frames, and 2) identity embedding from ArcFace [12] between the source images and the generated frames to evaluate the fidelity of the generated results. Apart from the quantitative metrics, we also conduct user studies to perceptually evaluate the result of different methods.